Question 16

A Pew Poll in 2010, found that 50% of adults aged 25-29 had access to only a cell phone, while 40% had access to both a cellphone and landline and 10% had access to only a landline. We wish to conduct a Goodness of Fit Test to see if the results of a poll of a random sample of 130 adults aged 25-29 are significantly different from the 2010 poll results. The table below presents the observed results.

Cell only Landline Only Both
Observed 69 39 22
Expected

Compute the expected count for the "BOTH" cell.
Answer #1

Given N = Sample size = 130

Proportion of adults having access to only a cell phone = 50% = 0.5

Proportion of adults having access a cellphone and landline = 40% = 0.4

Proportion of adults having access to only a landline = 10% = 0.1

Expected count for a cell = N * Proportion for the cell

Expected count for Cell only = 130 * 0.5 = 65

Expected count for Landline only cell = 130 * 0.1 = 13

Expected count for Both cell = 130 * 0.4 = 52